OpenStudy (anonymous):

find the indicated outputs for f(x)=4x^.. f(0) f(-1) f(2)

OpenStudy (anonymous):

what's this

OpenStudy (anonymous):

its one of my math questions

OpenStudy (anonymous):

just plug the values of x equal to 0 -1 and 2 turn by turn in given function and find the value for example ina first case x=0 f(0)=4*0=0

OpenStudy (anonymous):

let me retype it

OpenStudy (anonymous):

f(0)= f(-1)= f(2)= f(x)=4x^-4x

OpenStudy (anonymous):

tena dear u know Gauss(-Jordan) method ?

OpenStudy (anonymous):

Gauss(-Jordan) method ?

OpenStudy (anonymous):

f(0)=4*0-4*0=0 f(-1)=4*-1 -4*-1=-4+4=0 f(2)=4*2-4*2=0

OpenStudy (anonymous):

output of functon

OpenStudy (anonymous):

im lsot lol

OpenStudy (anonymous):

what happened tena32?

OpenStudy (anonymous):

u from ? Pak

OpenStudy (anonymous):

PAK

OpenStudy (anonymous):

KARACHI

OpenStudy (anonymous):

leave this tena profile and use other stuff